Find the square roots of z=5−12i
Found 2 solutions by abdulaleem, ikleyn:
Answer by abdulaleem(11)

Suppose that a+bi is a square root of 5 + 12i.
Then, (a+bi)^2 = (a^2 - b^2) + (2ab)i = 5 + 12i.
Equate real and imaginary parts:
a^2 - b^2 = 5
2ab = 12 ==> b = 6/a.
So, a^2 - (6/a)^2 = 5
==> a^2 - 36/a^2 = 5
==> a^4 -5a^2 - 36 = 0.
==> (a^2 -9)(a^2 + 4) = 0.
Since a must be real, a = 3 or -3.
This gives b = 2 or -2, respectively.
Thus, we have two square roots: 3+2i or -3-2i.

Answer by ikleyn(52781)

.
Suppose that a+bi is a square root of 5 - 12i. 

Then, (a+bi)^2 = (a^2 - b^2) + (2ab)i = 5 - 12i. 



Equate real and imaginary parts: 

a^2 - b^2 = 5 

2ab = -12 ==> b = -6/a. 


So, a^2 - (-6/a)^2 = 5 

==> a^2 - 36/a^2 = 5 

==> a^4 -5a^2 - 36 = 0. 

==> (a^2 -9)(a^2 + 4) = 0. 

Since a must be real, a = 3 or -3. 

This gives b = 2 or -2, respectively. 



Thus, we have two square roots: 3-2i or -3+2i.

--------------
Be aware: the final answer by the other tutor is incorrect !

===============
On complex numbers, there are the lessons in this site
    - Complex numbers and arithmetical operations on them
    - Complex plane
    - Addition and subtraction of complex numbers in complex plane
    - Multiplication and division of complex numbers in complex plane
    - Raising a complex number to an integer power
    - How to take a root of a complex number
    - Solution of the quadratic equation with real coefficients on complex domain
    - How to take a square root of a complex number
    - Solution of the quadratic equation with complex coefficients on complex domain

    - Solved problems on taking roots of complex numbers
    - Solved problems on arithmetic operations on complex numbers
    - Solved problem on taking square root of complex number (*)
    - Miscellaneous problems on complex numbers
    - Advanced problem on complex numbers
    - A curious example of an equation in complex numbers which HAS NO a solution


            ------>>>   Notice that your problem is  Problem 1  of the lesson marked  (*)  in the list.   <<<------


Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic  "Complex numbers".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.